OpenStudy (anonymous):

x^3/e^x lim as x ->infinity

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ x ^3}{ e^x }\]

OpenStudy (gorv):

\[a=\frac{ 1 }{ \frac{ 1 }{ a } }\]

OpenStudy (gorv):

so u can do the same

OpenStudy (gorv):

\[\lim_{x \rightarrow \infty}\frac{ \frac{ 1 }{ e^x} }{ \frac{ 1 }{ x^3} }\]

OpenStudy (gorv):

now it is 0/0 form apply L hopital